# Coarse-Grained Boltzmann Generators

### v1

@@ G.2 Ablation of Weight Clipping
As shown in Tab. 6, reweighting without clipping leads to large JS divergences and PMF errors, together with near-zero ESS, indicating severe weight degeneracy. In contrast, weight clipping restores stable estimates and substantially improves all metrics.
@@
Table 6: Quantitative comparison for alanine dipeptide across CG resolutions and different weight clipping thresholds, showing effects on JS, PMF errors, and ESS.
@@
Heavy Atom Proposal $0.0074\pm 0.0001$ $0.6051\pm 0.0102$
Heavy Atom Reweighted Proposal ($1\%$ clip) $0.0075\pm 0.0001$ $0.2493\pm 0.0126$ $0.2325\pm 0.0006$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.2364\pm 0.0584$ $8.3338\pm 3.2675$ $0.0000\pm 0.0001$
Core Beta Proposal 0.0072 $\pm$ 0.0001 $0.5236\pm 0.0099$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0079\pm 0.0001$ $0.2519\pm 0.0136$ 0.2510 $\pm$ 0.0006
Core Beta Reweighted Proposal ($0\%$ clip) $0.2398\pm 0.0923$ $10.0688\pm 5.6594$ $0.0001\pm 0.0001$
@@
Heavy Atom Proposal $0.0505\pm 0.0002$ $2.5861\pm 0.0353$
Heavy Atom Reweighted Proposal ($1\%$ clip) $0.0090\pm 0.0002$ $0.2758\pm 0.0109$ $0.1936\pm 0.0005$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.1565\pm 0.0399$ $6.0221\pm 2.5237$ $0.0001\pm 0.0001$
Core Beta Proposal $0.0540\pm 0.0002$ $3.7156\pm 0.0393$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0081\pm 0.0001$ 0.2325 $\pm$ 0.0093 $0.2239\pm 0.0005$
Core Beta Reweighted Proposal ($0\%$ clip) $0.5573\pm 0.1692$ $261.6973\pm 67.6421$ N/A
Figure 12: Metrics as a function of weight clipping ratio. (a) JS divergence. (b) PMF error. (c) ESS.
G.3 Free Energy of $\psi$ Dihedral
@@
Figure 13: Comparison of $\psi$ dihedral free energy profiles across training settings. All panels show the MD reference alongside CG-BG proposals before and after reweighting. Results correspond to: (a) Flow trained on unbiased data (Core Beta); (b) Flow trained on WT-MetaD data (Core Beta); (c) Flow trained on unbiased (Heavy Atom); (d) Flow trained on WT-MetaD data (Heavy Atom).
G.4 Bond Length
@@
Figure 14: Comparison of C–N bond length distributions across training settings. All panels show the MD reference alongside CG-BG proposals before and after reweighting. Results correspond to: (a) Flow trained on unbiased data (Core Beta); (b) Flow trained on WT-MetaD data (Core Beta); (c) Flow trained on unbiased data (Heavy Atom); (d) Flow trained on WT-MetaD data (Heavy Atom).

### v2

@@ G.4 Ablation of Weight Clipping
As shown in Tab. 7, Tab. 8 and Tab. 9, reweighting without clipping leads to large JS divergences and PMF errors, together with ESS, indicating severe weight degeneracy. In contrast, weight clipping restores stable estimates and substantially improves all metrics.
@@
Table 7: Quantitative comparison for alanine dipeptide across CG resolutions and different weight clipping thresholds. Atomistic BG results from previous studies (Tan et al., 2025a) are included for reference, where models are trained on implicit solvent datasets and evaluated after reweighting using the implicit solvent energy function. Reported values are computed against the explicit solvent MD reference. For alanine dipeptide, TarFlow is trained on biased simulation data, which leads to low ESS.
@@
Heavy Atom Proposal 0.0048(1) $0.2749(72)$
Heavy Atom Reweighted Proposal ($1\%$ clip) 0.0048(1) 0.2005(63) $0.5112(4)$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.0065(1)$ $0.2548(91)$ $0.4038(28)$
Core Beta Proposal $0.0058(1)$ $0.3662(85)$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0052(1)$ $0.2210(65)$ 0.5528(4)
Core Beta Reweighted Proposal ($0\%$ clip) $0.0065(1)$ $0.2833(84)$ $0.4592(50)$
@@
Heavy Atom Proposal $0.0510(3)$ $2.5527(382)$
Heavy Atom Reweighted Proposal ($1\%$ clip) $0.0063(1)$ $0.2277(66)$ $0.4115(4)$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.0058(1)$ $0.2163(61)$ $0.3603(23)$
Core Beta Proposal $0.0437(2)$ $1.7185(294)$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0057(1)$ $0.2093(58)$ $0.4818(4)$
Core Beta Reweighted Proposal ($0\%$ clip) $0.0058(1)$ $0.2015(55)$ $0.3939(11)$
Atomistic BGs Trained on Implicit Solvent Dataset
TarFlow $0.0548(2)$ $1.2696(314)$ $0.0034(18)$
ECNF++ $0.1451(19)$ $21.4005(4594)$ $0.2445(156)$
Table 8: Quantitative comparison for alanine tripeptide across CG resolutions and different weight clipping thresholds.
Model JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$)
Flow Trained on Unbiased Dataset
Heavy Atom Proposal $0.0065(1)$ $0.4071(73)$
Heavy Atom Reweighted Proposal ($1\%$ clip) 0.0056(1) 0.1957(52) $0.3201(4)$
Heavy Atom Reweighted Proposal ($0\%$ clip) $0.0110(1)$ $0.2926(184)$ $0.0368(192)$
Core Beta Proposal $0.0061(1)$ $0.3800(71)$
Core Beta Reweighted Proposal ($1\%$ clip) $0.0060(1)$ $0.2112(51)$ 0.4212(5)
Core Beta Reweighted Proposal ($0\%$ clip) $0.0414(2)$ $0.8451(142)$ $0.0018(10)$
Table 9: Quantitative comparison for alanine hexapeptide across different weight clipping thresholds.
Model JS ($\downarrow$) PMF ($\downarrow$) ESS ($\uparrow$)
Flow Trained on Unbiased Dataset
Core Beta Proposal $0.0134(1)$ $0.9801(130)$
Core Beta Reweighted Proposal ($1\%$ clip) 0.0100(1) 0.3646(81) 0.1231(3)
Core Beta Reweighted Proposal ($0\%$ clip) $0.1972(3)$ $8.0359(879)$ $0.0002(3)$
Figure 10: Metrics as a function of weight clipping ratio for flow models trained on unbiased datasets across different CG mappings and systems. (a) JS divergence, (b) PMF error, and (c) ESS after reweighting. The red dashed line indicates the 1% clipping ratio used throughout our experiments.
G.5 Free Energy of $\psi$ Dihedral
@@
Figure 11: Comparison of $\psi$ dihedral free energy profiles across training settings. All panels show MD reference distributions and CG-BG proposals before and after reweighting. (a) Dipeptide Core Beta (unbiased), (b) Dipeptide Core Beta (WT-MetaD), (c) Dipeptide Heavy Atom (unbiased), (d) Dipeptide Heavy Atom (WT-MetaD), (e) Tripeptide Core Beta, (f) Tripeptide Heavy Atom, (g) Hexapeptide Core Beta.
G.6 Bond Length
@@
Figure 12: Comparison of C–N bond length distributions across training settings. All panels show MD reference distributions and CG-BG proposals before and after reweighting. (a) Dipeptide Core Beta (unbiased), (b) Dipeptide Core Beta (WT-MetaD), (c) Dipeptide Heavy Atom (unbiased), (d) Dipeptide Heavy Atom (WT-MetaD), (e) Tripeptide Core Beta, (f) Tripeptide Heavy Atom, (g) Hexapeptide Core Beta.